# Diorama: Unleashing Zero-shot Single-view 3D Scene Modeling

### v1

@@ Abstract
Reconstructing structured 3D scenes from RGB images using CAD objects unlocks efficient and compact scene representations that maintain compositionality and interactability. Existing works propose training-heavy methods relying on either expensive yet inaccurate real-world annotations or controllable yet monotonous synthetic data that do not generalize well to unseen objects or domains. We present Diorama, the first zero-shot open-world system that holistically models 3D scenes from single-view RGB observations without requiring end-to-end training or human annotations. We show the feasibility of our approach by decomposing the problem into subtasks and introduce robust, generalizable solutions to each: architecture reconstruction, 3D shape retrieval, object pose estimation, and scene layout optimization. We evaluate our system on both synthetic and real-world data to show we significantly outperform baselines from prior work. We also demonstrate generalization to internet images and the text-to-scene task.
Figure 1: We propose Diorama: a system for zero-shot single-view 3D scene modeling. Our system produces a holistic 3D scene model given a single input image, representing both architectural elements and objects in cluttered indoor scenes. We show three example applications of our system. The input image is shown in the top left in each case. Other images are renderings of the output 3D scene from different camera viewpoints. (a) 3D scene model from a synthetic image input. (b) 3D scene model from an internet image input. (c) Text-to-scene pipeline where the input image on the left is generated from a text prompt.
@@ 1 Introduction
Modeling a 3D scene from a single image input is a cornerstone task in computer vision that is critical in real-world applications such as autonomous driving, augmented/virtual reality, and robotic perception. Recent works tackling 3D perception as a reconstruction problem tend to produce imperfect surface-only 3D meshes that are incompatible with modern graphics pipelines or simulation platforms [64, 54, 9]. An alternative strategy leverages a database of CAD models to model a 3D scene in an analysis-by-synthesis manner [46, 31, 27] that enables convenient scene creation and editing. Unfortunately, the potential of this strategy is constrained by the limited availability of annotated image–CAD scene pairs and limited generalization to diverse real-world images.
Creating 3D scenes typically involves recruiting expert 3D designers to manually build and assemble the scenes. The resulting 3D scenes are often oversimplified [24] or of limited accessibility to the research community with only images available [104, 71]. More recent works address the synthetic data domain gap by aligning CAD annotations to real-world RGBD scans [4, 77, 55], RGB videos [60] or panoramas [89]. However, they either ignore architecture and scene structure or do not reach scales appropriate for training.
@@
Our system consists of two major components: 1) Open-world perception for holistic scene understanding of the input image, including object recognition and localization, depth and normal estimation, architecture reconstruction and scene graph generation. 2) CAD-based scene modeling for assembly of a clean and compact 3D scene representation through CAD model retrieval, 9-DoF pose estimation, and semantic-aware scene layout optimization. Different from other zero-shot modular scene generation systems focusing on texture stylization [34], object reconstruction [106] and language description [26, 3], our work focuses on open-world holistic perception of objects, architecture, and spatial relationships. There are also works proposing to create synthetic data from images for robotic manipulation. However, current efforts [51, 13, 2, 86, 109, 10, 41] either use strong construction heuristics or only handle tabletop setups. We advocate our system as a more general approach to output cluttered compositional scenes that can be rearranged and edited. In summary, our contributions are:
@@
  • We ensure flexibility and generalization to real-world data with a modular design, proposing better and more robust solutions to general architecture reconstruction, 3D shape retrieval, zero-shot 9D pose estimation, and scene layout optimization.
  • We benchmark on synthetic and real-world data, and demonstrate the usability of our approach on in-the-wild images and text-to-scene applications.
Figure 2: Illustration of our Diorama system pipeline. The input image is processed in the open-world perception component (in orange box) through object instance segmentation, depth and normal estimation, architecture reconstruction and LLM-powered scene graph generation. The CAD-based scene modeling component (green box) then assembles a compositional 3D scene by retrieving and posing objects from a database and optimizing the overall scene layout. Multiple plausible scene arrangement hypotheses are produced as outputs.
@@ 2 Related work
Open-world 3D perception. Visual foundation models enabled various aspects of open-world 3D perception, including 3D shape understanding [103, 108, 91, 92, 107, 50], part-level 3D shape recognition [56, 1], 3D semantic segmentation [67, 102, 8, 16, 39], 3D instance segmentation [78, 62, 36], 3D object detection [58, 96], neural field rendering [43, 68] and language grounding [40, 15]. Most works obtains semantic-rich representation for open-world language query by directly leveraging CLIP features or learning to align with CLIP features. Some works [32, 25] propose to learn tokenized 3D features along with LLMs. In contrast, our work directly models complex, cluttered 3D scene from a single-view image and unlike prior work does not rely on any 3D modality inputs such as point clouds.
@@
Single-view scene modeling. There is a long history of work on scene modeling through CAD object retrieval and alignment [38, 35, 46, 47, 31, 48, 27, 59, 49]. IM2CAD [38] breaks down scene modeling into three major modules: room layout estimation, object recognition and scene configuration, similarly to our system. However, IM2CAD cannot easily generalize to unseen objects and oversimplifies complex architectures and arrangements in real-world images. More recent work explores end-to-end trained methods [46, 47, 31, 48]. Mask2CAD [46] jointly learns object detection, 3D shape embedding and direct pose regression. Patch2CAD [47] introduces a more robust patch-based joint embedding space for patch-level 3D shape retrieval. ROCA [31] leverages differentiable Procrustes optimization for pose estimation, and SPARC [48] employs an iterative “render-and-compare” strategy. DiffCAD [27] achieves state-of-the-art performance with a combination of probabilistic diffusion models trained in a weakly-supervised manner on synthetic data. However, all existing methods do not model architecture and require significant amounts of annotated data for training. Our system is training-free, requires no 3D supervision, and leverages large foundation models pretrained on large corpus internet data.
@@
Single-view scene reconstruction. There is much work on single-view scene reconstruction [29, 37, 64, 54, 9, 101, 17]. Due to the limited availability of real-world data, prior works train on synthetic 3D shapes. This typically leads to limitations such as oversmoothed and incomplete meshes, discarding of fine-grained scene structure, and real to synthetic domain gaps. Recent work implements modular zero-shot scene reconstruction using inpainting and image-conditioned 3D shape generation for each object [106]. In contrast, we produce a holistic scene model including both architecture and objects, and fine-grained object relations in highly cluttered scenes not handled by this prior work.
@@
Generative scene synthesis. Another line of work to synthetic 3D scene modeling is through learning scene configuration priors from large-scale data [24]. Various neural network-based approaches have explored capturing scene layout priors using CNNs [82, 70], relation graphs [83], transformers [66, 85], and diffusion models [87, 80, 52]. These approaches are restricted by training data quality and scale, and generally only produce furniture layouts on a horizontal plane. Recent work proposes to generate layout specifications using LLMs to mitigate the training data sparsity issues [26, 79, 28, 97, 3]. However, these approaches struggle to accurately position objects as LLMs do not encode fine-grained 3D spatial knowledge. In contrast, our system directly perceives images and converts them to plausible 3D layouts respecting the provided input.
3 System
@@
We present Diorama, a zero-shot open-world system that reconstructs synthetic 3D scenes from monocular image inputs, without end-to-end training as required by prior work [31, 48, 27]. Our system has two major components: 1) Open-world perception for holistic scene understanding of the input image, including object recognition and localization, depth and normal estimation, architecture reconstruction and scene graph generation (Fig. 2, orange box). 2) CAD-based scene modeling for assembly of a clean and compact 3D scene representation through CAD model retrieval, 9-DoF pose estimation, and semantic-aware scene layout optimization (Fig. 2, green box). As a modular system, its performance can be naturally improved by replacing each individual component with a dedicated approach of better performance. We choose different components based on the principles of open vocabulary, category agnosticism, and robustness to non-exact object matches. The reconstructed scenes benefit from certain designs of Diorama, including scene graph for maintaining spatial relationships among objects, shape retrieval for multiple semantically similar arrangements, planar architecture for physically plausible supporting objects, and layout optimization for refining objects poses based on spatial relationships. We describe each module in the following sections, with additional details in the supplement.
@@ 3.1 Open-world Perception
[…]
@@
Planar architecture reconstruction. Reconstructing the surrounding architecture enables better physical plausibility of the scene arrangement, by ensuring objects are supported on surfaces and are not floating in the air. We represent architecture as 3D planes following [75, 72] to better capture general complex room layout in the real world. We introduce PlainRecon, a simple yet effective architecture reconstruction pipeline leveraging inpainting [99], monocular depth [33, 95] and normal estimation [33] that handles both synthetic and real-world images. PlainRecon consists of three steps (see Fig. 3): 1) segment objects and inpaint to produce an “empty room” showing the complete architecture; 2) extract a point cloud and normals for each architecture element using depth estimation; and 3) fit 3D planes to architecture elements using normal-based clustering. Given planar segments obtained from normals-based clustering, we fit a plane equation to each point cloud of an architectural element and compute bounds in each plane. See the supplement for more details.
@@ 3.2 CAD-based scene modeling
Multimodal shape retrieval. We retrieve CAD objects from a database combining several 3D datasets [7, 11, 24, 14, 44]. Unlike 3D reconstruction pipelines, retrieval-based modeling creates a clean and compact scene representation which can more easily be used in downstream applications without post-processing, at the cost of potentially mismatches in geometry and appearance between 3D shapes and observed 2D objects. In essence, Diorama focuses on semantic matching of the observed objects under the open-world setting and focus on realistic object arrangements instead of precise geometric and texture modeling. We use DuoDuoCLIP [50] to encode text, images and 3D shapes into a joint embedding space. Specifically, we use a hierarchical retrieval strategy where we first retrieve from a text query to ensure that all 3D shape candidates lie in the same category group, then we sort retrieved shapes using an image query to re-rank shape candidates based on object appearance and take the best-matching one as the output.
Zero-shot object pose estimation. There is rich research literature in model-based pose estimation for unseen objects [30, 63, 88, 53, 5]. However, in our task we tackle a more challenging problem of estimating the pose of novel objects for semantically-matching but not necessarily exact geometrically-matching CAD objects. Our approach is a zero-shot category-level object pose estimation method similar to Goodwin et al. [30], that demonstrates generalization across unseen categories and removes the requirement for pose-labeled real image training data.
Specifically, we take advantage of the semantically-rich visual representation of a vision transformer [18] (ViT) learned in a self-supervised manner on large-scale data to compute object pose from feature similarity and correspondence. We pre-compute $T$ multiview renderings and depth maps for each retrieved CAD model $s_{i}$. Each query object crop $I^{o_{i}}$ and set of multiview images $\{I^{s_{i,j}}\}$ is encoded into normalized patch features $\mathcal{F}(I)$ using this ViT. For each patch embedding from the query image, we construct a semantic correspondence to a patch embedding from each rendering view with the minimal cyclical distance [30, 63]. We then compute a correspondence score as cosine similarity between patch embeddings and take the top-K correspondences. The query image is matched with the most similar multiview rendering based on the maximal averaged correspondence-wise feature similarity to produce a coarse pose hypothesis.
Since we have point clouds for both query object and CAD model we can also obtain 3D correspondences. Ideally, straightforward application of the Umeyama algorithm [81] with RANSAC on a set of 3D correspondences $q_{k}$ and $p_{k}$, can solve for a rigid body transform: rotation $R$, translation $t$, and a uniform scaling parameter $s$. However, we empirically find that $t$ and $s$ can be significantly influenced by undesired depth estimation and large object occlusion and truncation in image that result in visually implausible scene (object with extreme size or position deviation). We therefore adopt a small scale prediction network $\mathcal{F}_{\text{scale}}$ from GigaPose [63] for more robust scale estimation. The approach is illustrated in Fig. 4.
@@
Scene composition and optimization. We now have all the ingredients to compose a 3D scene from a set of retrieved and coarsely posed CAD models. Errors in the object poses may lead to violations in overall scene plausibility. Examples include inter-penetrations, flying objects, or gaps to the supporting objects. To address this, we use a differentiable semantic-aware scene layout optimization procedure to refine the coarse object poses. We separate the entire procedure into different stages to alleviate the difficulty of adjusting the 9D poses of all objects simultaneously. For each object, we first define its oriented bounding box (OBB) and then identify potential contact and support surfaces along with contact and support directions to optimize spatial support relations with other objects. We describe the optimization stages along with their terms below and illustrate the stage-wise optimization in Fig. 5.
@@
Orientation. For a pair of objects $o_{i}$ and $o_{j}$, and their associated contact surface normal $\mathbf{n}_{i}^{c}$ and support surface normal $\mathbf{n}_{j}^{s}$, we define the term $e_{\text{align}}$ as the $L_{2}$ norm of difference between the two normals $e_{\text{align}}=\sum_{(i,j)\in G}||\mathbf{n}_{i}^{c}-\mathbf{n}_{j}^{s}||_{2}$. During pose refinement, we maintain the front direction of each object projected on the horizontal plane by penalizing large disturbances. Specifically, we define the term $e_{\text{sem}}$ as the $L_{2}$ norm of residuals added to the projected vector, $e_{\text{sem}}=\sum_{i\in V}||\mathbf{v}_{i}^{*}-\mathbf{v}_{i}||_{2}$ where $\mathbf{v}^{*}$ indicates the initial front vectors before optimization.
@@
Placement. To ensure the supported object $o_{i}$ is placed on the supporting object $o_{j}$ for physical plausibility, we define the term $e_{\text{place}}$ as the $L_{1}$ norm of the distance vector from the center point of the contact surface to the support surface $e_{\text{place}}=\sum_{(i,j)\in G}||D_{\text{contact center}\rightarrow\text{support surface}}||_{1}$ We also define a spatial relation term $e_{\text{rel}}$ that deliberately maintains the relative $L_{2}$ distance between objects based on proximity relationships present in the scene, $e_{\text{rel}}=\sum_{(i,j)\in G}||(\mathbf{c}_{i}^{*}-\mathbf{c}_{j}^{*})-(\mathbf{c}_{i}-\mathbf{c}_{j})||_{2}$.
@@
Space. Each object should occupy its own 3D space and not penetrate other objects. We measure the amount of collision among objects and architecture elements following Zhang et al. [101], using the Separating Axis Theorem (SAT) to calculate the collision between convex polygons (cuboids). We penalize the penetration depths between any two objects on three orthogonal axes, $e_{\text{col}}=\sum_{\{i,\;j\}\in V}\sum_{axis}d_{\text{penetration}}$. In particular, we handle the usually ignored but common case where objects are placed inside another object (e.g. books on a bookshelf) by defining a supporting volume for each object and an optimization term $e_{\text{vol}}$ that penalizes the summed protrusions of each object from its containing volume bounds.
Refinement. Lastly, we encourage all objects to stick to their supporting surface by running the placement stage again. Additionally, we establish an adherence relationship between each object on the floor and potential walls using a heuristic algorithm based on a distance threshold. We hence also optimize object position and rotation based on the adherence relationship.

### v2

# Diorama: Unleashing Zero-shot Single-view 3D Indoor Scene Modeling

@@ Abstract
Reconstructing structured 3D scenes from RGB images using CAD objects unlocks efficient and compact scene representations that maintain compositionality and interactability. Existing works propose training-heavy methods relying on either expensive yet inaccurate real-world annotations or controllable yet monotonous synthetic data that do not generalize well to unseen objects or domains. We present Diorama, the first zero-shot open-world system that holistically models 3D scenes from single-view RGB observations without requiring end-to-end training or human annotations. We show the feasibility of our approach by decomposing the problem into subtasks and introduce better solutions to each: architecture reconstruction, 3D shape retrieval, object pose estimation, and scene layout optimization. We evaluate our system on both synthetic and real-world data to show we significantly outperform baselines from prior work. We also demonstrate generalization to real-world internet images and the text-to-scene task.
Figure 1: We propose Diorama: a system for zero-shot single-view 3D scene modeling. Our system produces a holistic 3D scene model given a single input image, representing both architectural elements and objects in cluttered indoor scenes. We showcase three examples of our system where the input image is shown in the top left in each case. Other images are renderings of the output 3D scene from different camera viewpoints. (a) 3D scene model from a synthetic image input. (b) 3D scene model from a real-world internet image input. (c) Text-to-scene pipeline where the input image on the left is generated from a text prompt.
@@ 1 Introduction
Modeling a 3D scene from a single image input is a cornerstone task in computer vision that is critical in real-world applications such as autonomous driving, augmented/virtual reality, and robotic perception. Recent works tackling 3D perception as a reconstruction problem tend to produce imperfect surface-only 3D meshes that are incompatible with modern graphics pipelines or simulation platforms [57, 47, 8]. An alternative strategy leverages a database of CAD models to model a 3D scene in an analysis-by-synthesis manner [40, 27, 24] that enables convenient scene creation and editing. Such representation, even with mismatched exterior geometry and texture, can be easily applied for robot learning [11, 98], leveraging human-labeled annotations, including physical properties, articulation parameters and complete interior geometry.
Creating 3D scenes typically involves intensive manual work by 3D designers. The resulting 3D scenes are oversimplified [21] or of limited accessibility to the public with only images available [93, 64]. More recent works address the synthetic data domain gap by aligning CAD objects to real-world RGBD scans [4, 70, 48], RGB videos [53] or panoramas [80]. However, they either ignore architecture or do not reach the scale appropriate for training.
@@
Our system consists of two major components: 1) Open-world perception for holistic scene understanding of the input image, including object recognition and localization, depth and normal estimation, architecture reconstruction and scene graph generation. 2) CAD-based scene modeling for assembly of a clean and compact 3D scene through CAD model retrieval, 9-DoF pose estimation, and semantic-aware scene layout optimization. Different from other zero-shot modular scene generation systems focusing on texture stylization [30], object reconstruction [95] and language description [23, 3], our work focuses on open-world holistic perception of objects, architecture, and spatial relationships. There are also works proposing to create synthetic scenes from images for robotic manipulation. However, current efforts [45, 11, 2, 98, 9] either use strong construction heuristics or only handle tabletop setups. We advocate our system as a more general approach to output cluttered compositional scenes that can be rearranged and edited. In summary, our contributions are:
@@
  • We ensure flexibility and generalization to real-world data with a modular design, proposing better and more robust solutions to planar architecture reconstruction, 3D shape retrieval, zero-shot 9D pose estimation, and stage-wise scene layout optimization.
  • We benchmark on synthetic and real-world data, and demonstrate the usability of our approach on real-world internet images and text-to-scene applications.
Figure 2: Illustration of the Diorama pipeline. The input image is processed in the open-world perception component (in orange box) through object instance segmentation, depth and normal estimation, architecture reconstruction and LLM-powered scene graph generation. The CAD-based scene modeling component (green box) then assembles a compositional 3D scene by retrieving and posing objects from a database and optimizing the overall scene layout. Multiple plausible scene arrangement hypotheses are produced as outputs.
@@ 2 Related work
Open-world 3D perception. Visual foundation models enabled various aspects of open-world 3D perception, including 3D shape understanding [97, 82, 50, 44], part-level 3D shape recognition [49, 1], 3D semantic segmentation [60, 92, 7, 14, 35], 3D instance segmentation [71, 55, 32], 3D object detection [51, 86], neural field rendering [37, 61] and language grounding [36, 13]. Most works obtain semantic-rich representation for open-world language query by directly leveraging CLIP feature or learning to align with CLIP feature. Some works [28, 22] propose to learn tokenized 3D feature along with LLMs. In contrast, our work directly models a complex, cluttered 3D scene from a single-view image and unlike prior work does not rely on any 3D modality inputs such as point clouds.
@@
Single-view scene modeling. There is a long history of work on scene modeling through CAD object retrieval and alignment [34, 31, 40, 41, 27, 42, 24, 52, 43]. IM2CAD [34] breaks down the task into separate modules: room layout estimation, object recognition and scene configuration. However, IM2CAD cannot easily generalize to unseen objects and oversimplifies complex architecture and arrangements in real-world images. More recent works explore end-to-end trained methods [40, 41, 27, 42]. Mask2CAD [40] jointly learns object detection, 3D shape embedding and pose regression. Patch2CAD [41] introduces a more robust patch-based joint embedding space for 3D shape retrieval. ROCA [27] leverages differentiable Procrustes optimization for pose estimation, and SPARC [42] employs an iterative “render-and-compare” strategy. DiffCAD [24] achieves state-of-the-art performance with a combination of probabilistic diffusion models trained in a weakly-supervised manner on synthetic data. However, all existing methods do not model architecture and require significant amounts of annotated data for training. Our system is training-free, requires no 3D supervision, and leverages large foundation models pretrained on large corpus data such that it generalizes to real-world images.
@@
Single-view scene reconstruction. There is much work on single-view scene reconstruction [33, 57, 47, 8, 91, 15]. Due to the limited availability of real-world data, prior works train on synthetic 3D shapes. This typically leads to limitations such as oversmoothed and incomplete meshes, discarding of fine-grained scene structure, and real to synthetic domain gaps. Recent work implements modular zero-shot scene reconstruction using inpainting and image-conditioned 3D shape generation for each object [95]. In contrast, we produce a holistic scene model including both architecture and objects, and fine-grained object relations in highly cluttered scenes not handled by this prior work.
@@
Generative scene synthesis. Another line of work to synthetic 3D scene modeling is through learning scene configuration priors from large-scale data [21]. Various neural network-based approaches have explored capturing scene layout priors using CNNs [75, 63], relation graphs [76], transformers [59, 78], and diffusion models [79, 73, 46]. These approaches are restricted by training data quality and scale, and generally only produce furniture layouts on a horizontal plane. Recent works [23, 72, 25, 87, 3] propose to generate layout specifications using LLMs to mitigate the training data sparsity issues. However, these approaches struggle to accurately position objects as LLMs do not encode fine-grained 3D spatial knowledge. In contrast, our system directly perceives images and converts them to plausible 3D layouts respecting the provided input.
3 Approach
@@
We present Diorama, a zero-shot open-world system that reconstructs synthetic 3D scenes from monocular image inputs, without end-to-end training as required by prior work [27, 42, 24]. Our system has two major components: 1) Open-world perception for holistic scene understanding of the input image, including object recognition and localization, depth and normal estimation, architecture reconstruction and scene graph generation (Fig. 2, orange box). 2) CAD-based scene modeling for assembly of a clean and compact 3D scene representation through CAD model retrieval, 9-DoF pose estimation, and semantic-aware scene layout optimization (Fig. 2, green box). Our system design follows the principles of open vocabulary, category agnosticism, and robustness to non-exact object matches. The modeled scenes benefit from certain designs of Diorama, including scene graph for maintaining spatial relationships among objects, shape retrieval for multiple semantically similar arrangements, planar architecture for physically plausible supporting objects, and layout optimization for refining objects poses based on spatial relationships. We describe each module in the following sections, with additional details in the supplement.
@@ 3.1 Open-world Perception
LMM-powered scene graph generation. To model complex real-world scenes, it is essential to understand the spatial relationships between objects. In particular, it is key to model the support hierarchy, which describes a complete static scene arrangement from distinct objects. We follow SoM [84] to leverage the visual grounding capability of Large Multimodal Model (GPT-4o). See the prompt example and details in the supplement. The output scene graph is formalized as $G=\langle V,E\rangle$, where each vertex $v_{i}\in V$ represents an object instance $o_{i}$ in the image along with its language caption and each edge $e=\langle v_{i},v_{j}\rangle\in E$ indicates $o_{i}$ is supported by $o_{j}$.
@@
Planar architecture reconstruction. Reconstructing the surrounding architecture enables better physical plausibility of the scene arrangement, by ensuring objects are supported on surfaces and are not floating in the air. We represent architecture as 3D planes following [68, 65] to better capture general complex room layout in the real world. We introduce PlainRecon, a simple yet effective architecture reconstruction pipeline leveraging inpainting [89], monocular depth and normal estimation [29] that handles both synthetic and real-world images. PlainRecon consists of three steps (see Fig. 3): 1) segment objects and inpaint to produce an “empty room” showing the complete architecture; 2) extract a point cloud and normals for each architecture element using depth estimation; and 3) fit 3D planes to architecture elements using normal-based clustering. Given planar segments obtained from normals-based clustering, we fit a plane equation to each point cloud of an architectural element and compute bounds in each plane. See the supplement for more technical details.
@@ 3.2 CAD-based scene modeling
Multimodal shape retrieval. We retrieve CAD objects from a combination of several 3D datasets [6, 12, 38]. Unlike reconstruction, retrieval-based modeling efficiently produces a clean and compact scene representation for downstream use without post-processing, but may introduce mismatches in geometry and appearance. In essence, Diorama focuses on efficient semantic matching of the observed objects under the open-world setting and understanding realistic object arrangements. We use DuoDuoCLIP [44] to encode text, images and 3D shapes into a joint embedding space. Specifically, we propose a hierarchical retrieval strategy where we first retrieve from a text query to ensure that all 3D shape candidates lie in the same category group, then we sort retrieved shapes using an image query to re-rank shape candidates based on object appearance and take the best-matching one as the output.
Zero-shot object pose estimation. We estimate the 9D pose of novel objects for semantically-matching but not necessarily exact geometrically-matching CAD objects. Our approach is a zero-shot model-based object pose estimation method that demonstrates generalization across unseen categories and removes the requirement for pose-labeled real image training data. We use the semantic-rich representation of visual foundation model (e.g. DinoV2 [58]) to compute 2D correspondences for a pair of query image and a multiview rendering of a CAD object. We compute correspondence scores as cosine similarity between patch embeddings and take the top-K correspondences. A coarse pose hypothesis is produced by taking the most similar multiview rendering with the maximal averaged correspondence-wise similarity. We then lift from 2D to 3D correspondences using depth. We apply the Umeyama algorithm [74] with RANSAC to solve for a rigid body transformation. We empirically find that the translation $t$ and the uniform scaling $s$ can be significantly influenced by undesired depth estimation and large object occlusion in image that result in visually implausible scene (object with extreme size or position deviation). We therefore adopt a small scale prediction network $\mathcal{F}_{\text{scale}}$ from GigaPose [56] for more robust scale estimation. The approach is illustrated in Fig. 4. See the supplement for more details.
@@
Stage-wise scene optimization. Errors in the object pose estimation may lead to violations in overall scene plausibility. Examples include inter-penetrations, flying objects, or gaps to the supporting objects. To address this, we use a differentiable semantic-aware scene layout optimization procedure to refine the coarse object poses. We separate the entire procedure into different stages to alleviate the difficulty of adjusting the 9D poses of all objects simultaneously. For each object, we first define its oriented bounding box (OBB) and then identify potential contact and support surfaces along with contact and support directions to optimize spatial support relations with other objects. We describe the optimization stages below and illustrate the stage-wise optimization in Fig. 5. See the supplement for details.
@@
Stage 1: Orientation. For a pair of objects $o_{i}$ and $o_{j}$, and their associated contact surface normal $\mathbf{n}_{i}^{c}$ and support surface normal $\mathbf{n}_{j}^{s}$, we define the term $e_{\text{align}}$ as the $L_{2}$ norm of difference between the two normals $e_{\text{align}}=\sum_{(i,j)\in G}||\mathbf{n}_{i}^{c}-\mathbf{n}_{j}^{s}||_{2}$. During pose refinement, we maintain the front direction of each object projected on the horizontal plane by penalizing large disturbances. Specifically, we define the term $e_{\text{sem}}$ as the $L_{2}$ norm of residuals added to the projected vector, $e_{\text{sem}}=\sum_{i\in V}||\mathbf{v}_{i}^{*}-\mathbf{v}_{i}||_{2}$ where $\mathbf{v}^{*}$ indicates the initial front vectors before optimization.
@@
Stage 2: Placement. To ensure the supported object $o_{i}$ is placed on the supporting object $o_{j}$ for physical plausibility, we define the term $e_{\text{place}}$ as the $L_{1}$ norm of the distance vector from the center point of the contact surface to the support surface $e_{\text{place}}=\sum_{(i,j)\in G}||D_{\text{contact center}\rightarrow\text{support surface}}||_{1}$ We also define a spatial relation term $e_{\text{rel}}$ that deliberately maintains the relative $L_{2}$ distance between objects based on proximity relationships present in the scene, $e_{\text{rel}}=\sum_{(i,j)\in G}||(\mathbf{c}_{i}^{*}-\mathbf{c}_{j}^{*})-(\mathbf{c}_{i}-\mathbf{c}_{j})||_{2}$.
@@
Stage 3: Space. Each object should occupy its own 3D space and not penetrate other objects. In particular, we handle the usually ignored but common case where objects are placed inside another object (e.g. books on a bookshelf) by defining a supporting volume for each object and an optimization term $e_{\text{vol}}$ that penalizes the summed protrusions of each object from its containing volume bounds.
Stage 4: Refinement. Lastly, we encourage all objects to stick to their supporting surface by running the placement stage again. We also measure the amount of collision among objects and architecture elements following Zhang et al. [91], using the Separating Axis Theorem (SAT) to calculate the collision between convex polygons (cuboids). We penalize the penetration depths between any two objects on three orthogonal axes, $e_{\text{col}}=\sum_{\{i,\;j\}\in V}\sum_{axis}d_{\text{penetration}}$.